OpenStudy (anonymous):

what must you add to complete the square? x^2+10x=8

OpenStudy (anonymous):

25-8=13

OpenStudy (anonymous):

oh wait. question probably means you want to write the lhs as a perfect square. half of 10 is 5 and \[5^2=25\] so add 25 to both sides .. sorry

OpenStudy (anonymous):

see basically the logic is to bring it to the form (x+b)^2=x^2+b2+2xb, so for getting this factorisation, see appropriately what to add. observe that the x term has co-efficient 2b and so if u hav a b^2 term u will form the complete square by the formula rite? so u add b^2

OpenStudy (anonymous):

if you add 25 to both sides you get \[x^2+10x+25=33\]or \[(x+5)^2=33\]